# Using Large Language Models to Tackle Fundamental Challenges in Graph Learning: A Comprehensive Survey

### v1

@@ Abstract
Graphs are a widely used paradigm for representing non-Euclidean data, with applications ranging from social network analysis to biomolecular prediction. Conventional graph learning approaches typically rely on fixed structural assumptions or fully observed data, limiting their effectiveness in more complex, noisy, or evolving settings. Consequently, real-world graph data often violates the assumptions of traditional graph learning methods, in particular, it leads to four fundamental challenges: (1) Incompleteness, real-world graphs have missing nodes, edges, or attributes; (2) Imbalance, the distribution of the labels of nodes or edges and their structures for real-world graphs are highly skewed; (3) Cross-domain Heterogeneity, graphs from different domains exhibit incompatible feature spaces or structural patterns; and (4) Dynamic Instability, graphs evolve over time in unpredictable ways. Recent advances in Large Language Models (LLMs) offer the potential to tackle these challenges by leveraging rich semantic reasoning and external knowledge. This survey provides a comprehensive review of how LLMs can be integrated with graph learning to address the aforementioned challenges. For each challenge, we review both traditional solutions and modern LLM-driven approaches, highlighting how LLMs contribute unique advantages. Finally, we discuss open research questions and promising future directions in this emerging interdisciplinary field. To support further exploration, we have curated a repository of recent advances on graph learning challenges: https://github.com/limengran98/Awesome-Literature-Graph-Learning-Challenges.
@@ 1 Introduction
Graphs provide a flexible and expressive framework for modeling complex systems composed of interconnected entities. Formally, a graph $G=(V,E,X_{V},X_{E})$ consists of a set of nodes $V$, edges $E$, node attributes $X_{V}$, and edge attributes $X_{E}$, where both nodes and edges may carry heterogeneous features. Unlike grid-structured data such as images or sequences, graphs encode non-Euclidean topologies with irregular neighborhood structures and varying connectivity patterns. To support learning on such structures, graph learning aims to derive low-dimensional node or graph-level representations $Z\in\mathbb{R}^{d}$ that preserve both structural dependencies and semantic information from the original graph [1, 2, 3]. Due to these representational advantages, graph learning has been successfully applied across a wide range of domains, including social network analysis [4], personalized recommendation [5], transportation optimization [6], financial modeling [7], and bioinformatics [8]. As a result, it has become a foundational technique for extracting knowledge from relational data and enabling models to reason over structured domains.
Despite this progress, real-world graph learning faces significant challenges that diverge from idealized assumptions often made in academic benchmarks. For instance, in social networks, cold-start users may lack sufficient links or profile features, making it difficult to learn reliable representations [5]. In fraud detection, the data is often highly imbalanced, where fraudulent behavior accounts for only a small fraction of all activities, and labels are sparse and noisy [9]. In smart cities, traffic systems involve heterogeneous and dynamically evolving road networks, posing difficulties for models to adapt to changing topologies in real time [6, 10, 11]. As summarized in Figure 1, these concrete scenarios highlight four fundamental challenges in real-world graph learning:
@@
A variety of traditional graph learning methods have been developed to address these challenges, achieving notable success in specific contexts. For instance, graph completion algorithms [27] and graph autoencoders [12] have been employed to handle missing data; techniques like GraphSMOTE [28] and re-weighted loss functions [29] are commonly used to mitigate data imbalance; and graph domain adaptation methods, including adversarial regularization [20, 30] and distribution alignment [23, 31], have been explored to address cross-domain heterogeneity. Additionally, dynamic graph models such as EvolveGCN [32] and DyRep [33] have been proposed to model evolving graph structures. While these methods are effective within their respective scopes, they often rely on task-specific designs and require significant domain expertise.
@@
In contrast, the emergence of large language models (LLMs) has opened up new opportunities for addressing the fundamental challenges of graph learning in a more unified and flexible way [34, 35, 36]. With their strong representational capacity, contextual reasoning abilities, and generalization potential, LLMs can extract rich semantic patterns from heterogeneous and noisy data, offering a promising complement—or even alternative—to traditional graph-specific approaches. To reflect this growing interest, Figure 2 shows publication trends over the past decade (up to May 2025) related to “Graph Learning” and “LLM-based Graph Learning”. The statistics indicate that not only has graph learning become an increasingly important topic, but the proportion of research exploring the integration of LLMs into graph learning has also been steadily rising.
@@
Figure 2: Annual publication trends in Graph Learning from January 2015 to May 2025, with a highlighted subset of works related to LLMs. The blue bars represent the total number of publications in Graph Learning each year, while the overlaid pink segments indicate the number of publications that are specifically related to LLMs. The percentages above each bar denote the proportion of LLM-related papers within the total Graph Learning publications for that year. Data sourced from Google Scholar (https://scholar.google.com).
@@
This survey provides a comprehensive review of current research at the intersection of graph learning and large language models, centered on the four fundamental challenges introduced above. To ensure broad coverage, we performed an extensive literature search across major databases (arXiv, ACM Digital Library, IEEE Xplore, Elsevier, and Springer, etc.) using keyword combinations such as “LLMs”, “graph learning”, “incompleteness”, “imbalance”, “heterogeneity”, “dynamic graphs”. We focus on recent publications (primarily 2015–2025) in top machine learning, data mining, and NLP venues, as well as relevant preprints. This survey includes papers that meet at least one of the following criteria: (1) the work explicitly addresses one of the four fundamental challenges in graph learning, or (2) the proposed solution incorporates LLMs as a central component. Based on these criteria, we initially collected and reviewed over 1,000 relevant papers. We then categorized them according to the type of challenge addressed and the strategies employed, ultimately selecting a curated set of over 260 representative works. For each challenge, we briefly discuss conventional (non-LLM) techniques to provide background context, while our primary focus is to highlight the novel contributions and advantages introduced by LLM-based approaches.
@@ 2 Overview of Graph Learning and LLMs
2.1 Graph Learning
@@
Random-Walk-Based Embedding Methods: Early approaches to graph learning draw from analogies to signal processing and dimensionality reduction. For example, graph signal processing (GSP) methods [38, 39] treat node features as signals on a graph and use spectral decomposition of the graph Laplacian to filter or extract features. Techniques like graph Fourier transforms and wavelets provide mathematical insight but often assume the graph is static and have high computational cost for large graphs. Another line of work uses matrix factorization, such as factorizing adjacency or Laplacian matrices [40] or methods like LINE [41] that factorize higher-order proximity matrices, to embed nodes in a low-dimensional space. These capture global graph structure but might miss local detail and again can be limited by data sparsity or missing entries.
A breakthrough in graph learning came with the introduction of random-walk and neural embedding methods. Algorithms like DeepWalk [42] and Node2Vec [43] perform truncated random walks on the graph to generate sequences of nodes, then apply techniques from word2vec to learn node embeddings where co-occurring nodes in walks have similar vectors. These methods, along with variants like metapath2vec [44] for heterogeneous graphs, are scalable and capture community structure, but they treat the graph as a static corpus of walks and do not incorporate node attribute features directly.
@@ 2.2 LLMs
LLMs are a class of deep neural networks pre-trained on massive corpora to model the probabilistic structure of natural language [59, 60]. By learning distributed semantic representations of tokens, LLMs are capable of capturing rich contextual dependencies, making them highly effective across a wide range of NLP tasks, such as text generation [61], question answering [62], machine translation [63], and sentiment analysis [64].
The evolution of LLMs represents a shift from manual feature engineering to end-to-end data-driven language understanding. Early NLP systems relied on symbolic rules and statistical models (e.g., regular expressions [65], HMMs [66], and CRFs [67]), which were limited by local context and domain specificity. The introduction of the Transformer architecture by Vaswani et al. [68] in 2017 enabled global attention-based modeling, laying the foundation for scalable deep language models.
Subsequent models such as BERT [69] introduced bidirectional context modeling through masked language modeling, while GPT-3 [70] demonstrated the power of the “Scaling Law” by achieving strong few-shot capabilities with 175 billion parameters. The introduction of ChatGPT [71], fine-tuned via reinforcement learning with human feedback (RLHF), brought conversational alignment and intent understanding to new levels of human-like interaction.
Recent models, including GPT-4 [72], Gemini [73], and DeepSeek-MoE [74], mark a new phase of innovation. These models integrate visual encoders, multimodal fusion, and mixture-of-experts (MoE) architectures, enabling cross-modal understanding while optimizing computation. DeepSeek-MoE, for instance, uses dynamic expert routing to reduce computational overhead without sacrificing accuracy, offering a scalable path to deployment.
LLMs, by virtue of their training, encapsulate a broad range of knowledge and have shown surprising abilities in reasoning, abstraction, and generalization. This makes them attractive for addressing graph challenges as follows:
Semantic reasoning: LLMs excel at inferring latent information and making logical connections based on context. They can understand descriptions of entities and predict relationships that are not explicitly stated but are likely given the context. For example, given a partial description of a node, an LLM might deduce other attributes or links that should exist.
Cross-modal and cross-domain integration: LLMs can serve as a bridge between different data modalities and domains. They can take textual input and integrate it with structural information by producing a unified representation or by generating structured outputs. This is highly useful for heterogeneous graphs – e.g., aligning a gene interaction network with a drug-side-effect network by reasoning over biomedical texts that link the two.
Generative augmentation: LLMs have the ability to generate content – whether it be new text, or by extension new graph elements described by that text. This generative power can create synthetic nodes or edges with plausible attributes for data augmentation, address class imbalance by generating additional samples, or even hypothesize new connections in knowledge graphs.
This background section has outlined the principles of graph learning and LLMs. Figure 3(b) illustrates representative paradigms for leveraging LLMs in graph learning, where LLMs interact with graph data in different roles to support structure enhancement, representation learning, and predictive reasoning. In the following section, we bring these together by examining how LLMs have been applied to each of the four fundamental graph learning challenges. We will see that for each challenge, LLMs introduce new ways to overcome the limitations of traditional graph algorithms, whether by injecting external knowledge, balancing data through generation, aligning heterogeneous information, or reasoning about changes.
3 Related Work and Current Techniques for Graph Learning Challenges
@@ 3.1 Incompleteness in Graphs
In the real world, graph data is often incomplete due to missing node attributes, edges or labels (as shown in Figure 1), which significantly impacts the accuracy and generalization ability of graph models. By leveraging LLMs for graph data completion and the inference of missing information, the performance of graph learning models can be significantly improved, even in scenarios with incomplete or missing information. We categorize current approaches to incomplete graphs into three major directions (as shown in Figure 4): (i) Robust Graph Learning, which focuses on making models resilient to missing data; (ii) Few-Shot Graph Learning, which addresses scenarios with extremely limited labeled data or structure by transferring knowledge; and (iii) Knowledge Graph Completion, which is a prominent sub-area dealing with inferring missing facts in large knowledge graphs. The relevant references and categorization are presented in Table 1.
@@
Table 1: Literature Summary of LLM-based Methods for Handling Incompleteness in Graphs.
Domain Incomplete Type Reference (Method—Citation) Robust Graph Learning Node LLM4NG[75], LLM-TAG[76], SPLLM[77] Label LLMGNN[78] Mixed GraphLLM[79], PROLINK[80], UnIMP[81] Few-Shot Graph Learning Structure LinkGPT[82], AnomalyLLM[83] Mixed LLMDGCN[84], HGT[85] FlexKBQA[86], KGQG[87] Knowledge Graph Completion Node LLM-KGC[88], GS-KGC[89], GLTW[90] Label KGs-LLM[91] Mixed FSKG[92], KGLLM[93], KICGPT[94] RL-LLM[95], GoG[96], KoPA[97] LLMKG[98], DIFT[99] CP-KGC[100], MuKDC[101]

### v2

# A Survey of Large Language Models for Data Challenges in Graphs

@@ Abstract
Graphs are a widely used paradigm for representing non-Euclidean data, with applications ranging from social network analysis to biomolecular prediction. While graph learning has achieved remarkable progress, real-world graph data presents a number of challenges that significantly hinder the learning process. In this survey, we focus on four fundamental data-centric challenges: (1) Incompleteness, real-world graphs have missing nodes, edges, or attributes; (2) Imbalance, the distribution of the labels of nodes or edges and their structures for real-world graphs are highly skewed; (3) Cross-domain Heterogeneity, graphs from different domains exhibit incompatible feature spaces or structural patterns; and (4) Dynamic Instability, graphs evolve over time in unpredictable ways. Recently, Large Language Models (LLMs) offer the potential to tackle these challenges by leveraging rich semantic reasoning and external knowledge. This survey focuses on how LLMs can address four fundamental data-centric challenges in graph-structured data, thereby improving the effectiveness of graph learning. For each challenge, we review both traditional solutions and modern LLM-driven approaches, highlighting how LLMs contribute unique advantages. Finally, we discuss open research questions and promising future directions in this emerging interdisciplinary field. To support further exploration, we have curated a repository of recent advances on graph learning challenges: https://github.com/limengran98/Awesome-Literature-Graph-Learning-Challenges.
@@ 1 Introduction
Graphs provide a flexible and expressive framework for modeling complex systems composed of interconnected entities. Unlike grid-structured data such as images or sequences, graphs encode non-Euclidean topologies with irregular neighborhood structures and varying connectivity patterns. To support learning on such structures, graph learning aims to derive low-dimensional node or graph-level representations that preserve both structural dependencies and semantic information from the original graph [1, 2, 3, 4]. Due to these representational advantages, graph learning has been successfully applied across a wide range of domains, including social network analysis [5], personalized recommendation [6], transportation optimization [7], financial modeling [8], and bioinformatics [9, 10]. As a result, it has become a foundational technique for extracting knowledge from relational data and enabling models to reason over structured domains.
Despite this progress, real-world graph data often exhibits complexities, including sparsity, noise, domain divergence, and dynamic evolution, introducing practical difficulties for graph learning [11]. For instance, in social networks, cold-start users may lack sufficient links or profile features, making it difficult to learn reliable representations [6]. In fraud detection, the data is often highly imbalanced, where fraudulent behavior accounts for only a small fraction of all activities, and labels are sparse and noisy [12]. In smart cities, traffic systems involve heterogeneous and dynamically evolving road networks, posing difficulties for models to adapt to changing topologies in real time [7, 13, 14]. As summarized in Figure 1, these concrete scenarios highlight four fundamental data-centric challenges in real-world graphs:
@@
A variety of graph learning methods have been developed to address these challenges, achieving notable success in specific contexts. For instance, graph completion algorithms [30] and graph autoencoders [15] have been employed to handle missing data; techniques like GraphSMOTE [31] and re-weighted loss functions [32] are commonly used to mitigate data imbalance; and graph domain adaptation methods, including adversarial regularization [23, 33] and distribution alignment [26, 34], have been explored to address cross-domain heterogeneity. Additionally, dynamic graph models such as EvolveGCN [35] and DyRep [36] have been proposed to model evolving graph structures. While these methods are effective within their respective scopes, they often rely on task-specific designs and require significant domain expertise.
@@
In contrast, the emergence of large language models (LLMs) has opened up new opportunities for mitigating these fundamental challenges [37, 38, 39]. With their strong representational capacity, contextual reasoning abilities, and generalization potential, LLMs can extract rich semantic patterns from heterogeneous and noisy data. These capabilities enable LLMs to augment graph data by inferring missing information [40], synthesizing data for underrepresented classes [41], aligning heterogeneous attributes [42], and capturing temporal evolution or structural changes in dynamic graphs [43]. By addressing these issues, LLMs offer a promising complement to traditional graph learning approaches. To reflect this growing interest, Figure 2 shows publication trends over the past decade (up to July 2025) related to “Graph Learning” and “LLM-based Graph Learning”. The statistics indicate that not only has graph learning become an increasingly important topic, but the proportion of research exploring the integration of LLMs into graph learning has also been steadily rising.
@@
Figure 2: Annual publication trends in Graph Learning from January 2015 to July 2025, with a highlighted subset of works related to LLMs. The blue bars represent Graph Learning (non-LLM) publications and the pink segments represent LLM-related publications; the stacked height indicates the total per year. The percentages above each bar denote the proportion of LLM-related papers within the total for that year. Data sourced from Google Scholar (https://scholar.google.com).
@@
This survey provides a comprehensive review of current research at the intersection of graph learning and large language models, centered on the four fundamental challenges introduced above. To ensure broad coverage, we performed an extensive literature search across major databases (arXiv, ACM Digital Library, IEEE Xplore, Elsevier, and Springer, etc.) using keyword combinations such as “LLMs”, “graph learning”, “incompleteness”, “imbalance”, “heterogeneity”, “dynamic graphs”. We focus on recent publications (primarily 2015–2025) in top machine learning, data mining, and NLP venues, as well as relevant preprints. This survey includes papers that meet at least one of the following criteria: (1) the work explicitly addresses one of the four fundamental challenges in graph learning, or (2) the proposed solution incorporates LLMs as a central component. Based on these criteria, we initially collected and reviewed over 1,000 relevant papers. We then categorized them according to the type of challenge addressed and the strategies employed, ultimately selecting a curated set of over 380 representative works. For each challenge, we briefly discuss conventional (non-LLM) techniques to provide background context, while our primary focus is to highlight the novel contributions and advantages introduced by LLM-based approaches.
@@ 2 Overview of Graph Learning and LLMs
2.1 Graph Notation Definition
We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X}_{V},\mathbf{X}_{E})$, where $\mathcal{V}$ is the set of $n$ nodes, $\mathcal{E}$ is the set of edges, $\mathbf{X}_{V}\in\mathbb{R}^{|\mathcal{V}|\times d_{v}}$ is the node feature matrix, $\mathbf{X}_{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$ is the edge feature matrix, and $\mathbf{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, where $\mathbf{A}_{ij}=1$ if $(v_{i},v_{j})\in\mathcal{E}$. The degree matrix is $\mathbf{D}$ with $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. The normalized graph Laplacian is defined as: $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}.$
The types of graph data involved in this work can be broadly categorized as follows. Homogeneous graph refers to a graph in which all nodes and edges belong to the same type. Heterogeneous graph refers to a graph where nodes and/or edges belong to multiple types, such as user/item nodes or different edge relations. Attributed graph refers to a graph where nodes and/or edges are associated with real-valued features, i.e., $\mathbf{X}_{V}$ and/or $\mathbf{X}_{E}$ are non-empty. Knowledge graph refers to a directed, multi-relational graph represented as triplets $(h,r,o)$, where $h\in\mathcal{V}$ is the head entity, $o\in\mathcal{V}$ is the object entity, and $r\in\mathcal{R}$ is a relation type. Dynamic graph refers to a sequence of time-evolving graphs $\{\mathcal{G}_{t}\}_{t=1}^{T}$ in which the node set, edge set, or features may change over time.
2.2 Graph Learning Methods
@@
Traditional Graph Learning Methods: A wide range of traditional techniques have been developed to learn graph representations. Early methods such as spectral clustering [45] and LINE [46] learn embeddings by factorizing adjacency or similarity matrices, capturing global connectivity patterns. Random-walk-based approaches like DeepWalk [47], Node2Vec [48], and metapath2vec [49] model local contexts through node sequences generated by truncated walks, offering better scalability and flexibility. In parallel, autoencoder-based models aim to reconstruct graph structures from latent representations. Notable examples include Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) [50], which adopt graph convolutional encoders with deterministic or probabilistic decoders. Further developments such as Adversarially Regularized Graph Autoencoders (ARGA) [51] and Graph Attention Autoencoders (GATE) [52] introduce adversarial training and attention mechanisms to improve expressiveness and robustness. Despite their foundational role, traditional methods often fall short in handling the complexity of real-world graph data.
@@ 2.3 LLMs for Graph Learning
LLMs are a class of deep neural networks pretrained on massive text corpora to capture the statistical and semantic structure of natural language [67, 68]. By learning distributed token representations and contextual relationships, LLMs achieve impressive performance across a broad spectrum of NLP tasks, including text generation [69], question answering [70], and machine translation [71].
The development of LLMs reflects a paradigm shift from task-specific pipelines to general-purpose language understanding. Early NLP relied on symbolic or statistical models (e.g., HMMs [72], CRFs [73]), which were constrained by domain specificity and limited context. The introduction of the Transformer architecture [74] enabled global attention and deep sequence modeling, which laid the foundation for modern LLMs. Subsequent milestones such as BERT [75], GPT-3 [76], and ChatGPT [77] progressively advanced context modeling, few-shot reasoning, and conversational alignment. Recent systems like GPT-4 [78] and DeepSeek-MoE [79] further scale capacity and introduce modular expert routing, enabling more efficient and multimodal learning.
Despite being originally designed for natural language, LLMs exhibit emergent capabilities—such as abstract reasoning, flexible generalization, and heterogeneous input integration—that extend far beyond textual tasks. As shown in Figure 3(b), LLMs can complement graph learning by generating structural hypotheses [80], enriching semantic content [40], and aligning cross-domain knowledge [42]. Viewed through the lens of graph learning, these abilities make LLMs well-suited for enhancing data quality and robustness in graph-based applications:
Semantic reasoning: By leveraging rich contextual understanding, LLMs can infer latent or missing information from incomplete inputs. In the graph domain, this enables the imputation of missing node or edge attributes, the prediction of potential links, and the interpretation of sparsely connected substructures. Such capabilities directly address the challenge of incompleteness in real-world graphs, where data sparsity or noise is common [40].
Cross-domain and cross-modal alignment: LLMs are trained on diverse sources of knowledge and can serve as translators between incompatible domains or feature spaces. This makes them particularly useful in cross-domain heterogeneous graphs [42]—for example, aligning product graphs across e-commerce platforms that differ in attribute schemas, category taxonomies, and language conventions. LLMs can bridge such gaps by reasoning over textual descriptions and contextual cues to identify semantically equivalent items across domains.
Generative augmentation: As generative models, LLMs can synthesize new graph components—such as nodes, edges, or subgraphs—conditioned on textual or structural prompts. This ability can be used to create balanced training samples for underrepresented classes, thereby mitigating data imbalance [41], or to simulate hypothetical future graph states for handling dynamic instability [80].
In the following sections, we examine each of the four fundamental data-centric challenges in turn, analyzing how LLMs have been used to mitigate them, and what open questions remain at this intersection.
3 Current Techniques for Graph Challenges
@@ 3.1 Incompleteness in Graphs
In the real world, graph data is often incomplete due to missing node attributes, edges or labels, which significantly impacts the accuracy and generalization ability of graph models. By leveraging LLMs for graph data completion and the inference of missing information, the performance of graph learning models can be significantly improved, even in scenarios with incomplete or missing information. We categorize current approaches to incomplete graphs into three major directions (as shown in Figure 4): (1) Robust Graph Learning, which focuses on making models resilient to missing data; (2) Few-Shot Graph Learning, which addresses scenarios with extremely limited labeled data or structure by transferring knowledge; and (3) Knowledge Graph Completion, which is a prominent sub-area dealing with inferring missing facts in large knowledge graphs. The relevant references and categorization are presented in Table 1.
@@
Table 1: LLM-based methods for handling incompleteness in graphs, grouped by domain and incomplete type, with representative methods, datasets, metrics, and downstream tasks.
Domain Incompleteness Method Typical Datasets Common Metrics Downstream Tasks Robust Graph Learning Node LLM4NG [41] Cora [81], PubMed [82], ogbn-arxiv [83] Accuracy Node Classification LLM-TAG [84] Cora [81], Citeseer [85], PubMed [82], Arxiv-2023 [86] Accuracy Node Classification SPLLM [87] PeMS03 [88], PeMS04 [88], PeMS07 [88] MAE, RMSE, MAPE Spatiotemporal Forecasting Label LLMGNN [89] Cora [81], Citeseer [85], PubMed [82], ogbn-arxiv [83], ogbn-products [83], Wikics [90] Accuracy Node Classification Mixed GraphLLM [91] Synthetic Data [92] Exact Match Accuracy Graph Reasoning PROLINK [93] FB15k237 [94], Wikidata68K [95], NELL-995 [96] MRR, Hits@N Knowledge Graph Completion UnIMP [97] BG [98], ZO [98], PK [98], BK [98], CS [98], ST [98], PW [98], BY [98], RR [98], WM [98] RMSE, MAE Data Imputation Few-Shot Graph Learning Structure LinkGPT [99] AmazonSports [100], Amazon-Clothing [100], MAG-Geology [101], MAG-Math [101] MRR, Hits@N Link Prediction AnomalyLLM [102] UCI Messages [103], Blogcatalog [104], T-Finance [105], T-Social [105] AUC Anomaly Detection Mixed LLMDGCN [106] Cora [81], Citeseer [85], PubMed [82], Religion [107] Accuracy Node Classification HeGTa [108] IM-TQA [109], WCC [110], HiTab [111], WTQ [112], TabFact [113] Macro-F1, Accuracy Table Understanding FlexKBQA [114] GrailQA [115], WebQSP [116], KQA Pro [117] Exact Match, F1, Accuracy Knowledge Graph Question Answering KGQG [118] WebQuestions [119], PathQuestions [120] BLEU-4, ROUGE-L, Hits@N Knowledge Graph Question Answering Knowledge Graph Completion Node LLM-KGC [121] ILPC [122] MRR, Hits@N Knowledge Graph Completion GS-KGC [123] WN18RR [124], FB15k-237 [94], FB15k-237N [125], ICEWS14 [126], ICEWS05-15 [127] Hits@N Knowledge Graph Completion GLTW [128] FB15k-237 [94], WN18RR [124], Wikidata5M [129] MRR, Hits@N Link Prediction Label KGs-LLM [130] Wikipedia [130] F1, Precision, Recall Knowledge Graph Generation Mixed FSKG [40] WN18RR [124], FB15k-237 [94] MRR, Hits@N Knowledge Graph Completion KGLLM [131] WN11 [132], FB13 [132], WN18RR [124], YAGO3-10 [124] Accuracy, MRR, Hits@N Link Prediction, Knowledge Graph Completion KICGPT [133] FB15k-237 [94], WN18RR [124] MRR, Hits@N Link Prediction RL-LLM [134] Electronics, Instacart [134] Precision, Recall, Accuracy Knowledge Graph Completion GoG [135] Synthetic Data [135] Hits@N Knowledge Graph Question Answering KoPA [136] UMLS [137], CoDeX-S [125], FB15K-237N [125] F1, Precision, Recall, Accuracy Knowledge Graph Completion LLMKG [138] Templates Easy [138], Templates Hard [138] Strict Metrics, Flexible Metrics Knowledge Graph Completion DIFT [139] WN18RR [124], FB15k-237 [94] MRR, Hits@N Link Prediction, Knowledge Graph Completion CP-KGC [140] WN18RR [124], FB15k-237 [94], UMLS [137] MRR, Hits@N Knowledge Graph Completion MuKDC [141] NELL [142], Wiki [129] MRR, Hits@N Knowledge Graph Completion